OpenStudy (anonymous):

Calculus: What does it mean to take the derivative of something with respect to something? I know how I can compute the derivative of something, but what does it mean to take it with respect to something else?

OpenStudy (anonymous):

It means solve for the particular derivative

OpenStudy (amistre64):

how does this change when we change this other thing

OpenStudy (anonymous):

can you be more concrete?

OpenStudy (amistre64):

how fast is the height of a ball changing after 5 seconds of being dropped from 300 meters?

OpenStudy (anonymous):

taking the derivative with respect to x means finding the change in the x derivative or partial derivative

OpenStudy (amistre64):

a derivative is a rate of change; the "respect to" is relating it to some variable

OpenStudy (anonymous):

so like finding the equation for the tangent of a function at any point?

OpenStudy (anonymous):

slope of tangent*

OpenStudy (amistre64):

how long can you last in a boxing ring ... with respect to Mike Tyson? how long can you last in a boxing ring ... with respect to Gary Bussey?

OpenStudy (anonymous):

so you wouldn't solve for the equation but for the particular value

OpenStudy (amistre64):

the derivative gives you another equation, which can be used to determine the slope of the tangent (the measure of the rate of change) at a given value

OpenStudy (anonymous):

I know that

OpenStudy (anonymous):

But when you're taking it with respect to something, you're basically asking what the slope of that tangent is at that point, instead of the equation of the slope of the tangent at any point?

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

ok thanks!

OpenStudy (amistre64):

yw

OpenStudy (anonymous):

Hmm, let me ask another question, because I think I still might be understanding it wrongly: derivative of ln(x^2) with respect to x^2, according to khan academy, is 1/(x^2) But if you solve for the particular value, wouldn't that actually mean you have to find the derivative of ln( (x^2)^2 )?

OpenStudy (amistre64):

if we let: a = x^2 ln(a)' = 1/a

OpenStudy (amistre64):

if we want to know what it is wrt "x"; then let a(x) = x^2 ln(a)' = a'/a

OpenStudy (anonymous):

Well I see I'm interpreting it wrong, I actually have to find the derivative of ln(x^2), but then inside the derivative, substitute x^2 for x

OpenStudy (anonymous):

yet, derivative of ln(x^2) is actually (1/(x^2))*2x

OpenStudy (amistre64):

keep in mind that the chain rule is always at play.\[\frac d{dx}(ln(x^2))=\frac{\frac d{dx}(x^2)}{x^2}\to\ \frac{2x}{x^2}\]

OpenStudy (amistre64):

what would you say is the derivative of x^3 by chance?

OpenStudy (anonymous):

but you said if we let: a = x^2 ln(a)' = 1/a

OpenStudy (amistre64):

if we are deriving "with respect to" x^2, then yes ....

OpenStudy (anonymous):

but I guess the how "with respect to" is meant by khan academy isn't formally the same thing as how you defined it?

OpenStudy (anonymous):

Is it wrong to think of whatever you are taking respect to as one variable. So in your example if you are taking with respect to x^2 you need to think of it as "a", only one variable, instead of x^2....?

OpenStudy (amistre64):

my definitions tend to be based on my own comprehension

OpenStudy (anonymous):

Is there no formal definition of this?

OpenStudy (amistre64):

im sure there is, its prolly located in a dusty textbook on a shelf someplace :)

OpenStudy (anonymous):

x)

OpenStudy (anonymous):

I mean I kind of understand it both ways, both in your definition and how khan academy uses it, but that's not getting me anywhere, I need some definition that's always true

OpenStudy (amistre64):

the only formal definition of a derivative that I can recall is \[\lim_{\Delta x\to\ 0}\frac{f(x+\Delta x)-f(x)}{\Delta x}\]

OpenStudy (anonymous):

yeah, I know that one. But when people say we take derivative of something with respect to something they're just getting me confused because I don't know what they mean, they could mean so many different things...

OpenStudy (amistre64):

then lets try this: what would you say is the derivative of x^3 by chance?

OpenStudy (anonymous):

2x^3

OpenStudy (anonymous):

uhh sorry!

OpenStudy (amistre64):

ok, now why did you assume it was that?

OpenStudy (anonymous):

3x^2 Well, some rule.

OpenStudy (amistre64):

3x^2 ... of course :) but your missing something from it your missing the chain rule part

OpenStudy (amistre64):

do we know what this is "with respect to" ?

OpenStudy (anonymous):

well not really

OpenStudy (amistre64):

then we allow the chain rule to define the unknown the derivative of: x^3 is, 3x^2 * x'

OpenStudy (amistre64):

now, when i say that this is wrt x; what does x' equate to?

OpenStudy (anonymous):

wrt?

OpenStudy (amistre64):

w.ith r.espect t.o wrt

OpenStudy (anonymous):

3x^2 * 1x^(1-1) = 3x^2

OpenStudy (amistre64):

if i were to say, what is the derivative of x^3, wrt n; what would we get?

OpenStudy (anonymous):

3x^2 * n' ?

OpenStudy (amistre64):

not n': x' we assume by the construct of the question that x is a function of n; or rather x(n) Dn [x(n)^3] = 3(x(n)^2)*x'(n) which is usually cleaned up as Dn [x^3] = 3x^2 * x'

OpenStudy (amistre64):

when y is a function of x: y(x) and you take the derivative wrt x; do you end up with x'? No. the derivative of y wrt x is y' the derivative of x wrt n is x' the derivative of r wrt t is r' the "wrt" defines the independant variable that we are changing to see how it affect the dependant variable

OpenStudy (amistre64):

until we know the "wrt"; the chain rule just pops out a ' version :)

OpenStudy (amistre64):

if you have an example we can run thru, something that your having issues on; that might be more productive

OpenStudy (anonymous):

I'm not sure if I'm getting it but I'll reread it in a while, though you'll have to excuse me because I have to go afk for a while

OpenStudy (anonymous):

thanks for the help so far though

OpenStudy (amistre64):

ill be in class from 130 to 5 or so, so i might not get back to this till 2mrw; but im sure there are others capable of help you as well

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

I'm still not getting this wrt... Yes, I reread it

OpenStudy (anonymous):

Dn [x^3] = 3x^2 * x' What is x' here?

OpenStudy (anonymous):

And Dn?

OpenStudy (anonymous):

derivative with respect to n?

OpenStudy (turingtest):

that n makes no sense to me either

OpenStudy (anonymous):

oh glad I'm not alone

OpenStudy (turingtest):

usually you may see\[D_x[f(x)]=f'(x)\]which is the derivative with respect to x

OpenStudy (anonymous):

\[= \frac{dy}{dx}f(x) ?\]

OpenStudy (turingtest):

you could also see\[\frac{d^n}{dx^n}f(x)=f^{(n)}(x)\]which would mean the \(n^{th}\) derivative with respect to x

OpenStudy (turingtest):

what you wrote right now makes little sense

OpenStudy (turingtest):

usually, in single-variable calculus, y is just a variable used to represent a function of x f(x)

OpenStudy (turingtest):

so\[\frac{dy}{dx}=\frac d{dx}f(x)=f'(x)\]

OpenStudy (anonymous):

how do you read the second one?

OpenStudy (anonymous):

difference in ? over difference in x

OpenStudy (turingtest):

so what the heck is\[\frac{dy}{dx}f(x)=\frac d{dx}[f(x)]f(x)=f'(x)f(x)\]I guess... this is an ill- formed expression how to read it? all these things are read the derivative with respect to (whatever is bottom, d/dx is wrt x, d/dy is wrt y, d/d(purple) is the derivative with respect to purple, etc.)

OpenStudy (turingtest):

\[\frac {d^n}{dx^n}f(x)=f^{(n)}(x) \]is read as "the \(n^{th}\) derivative with respect to x of f of x"

OpenStudy (anonymous):

all right. But I still don't get this with respect to. Can you introduce me to a formal definition?

OpenStudy (turingtest):

formal... hm...

OpenStudy (anonymous):

But wait, shouldn't that last latex expression be read as "the nth derivative with respect to x^n of f(x)" ?

OpenStudy (turingtest):

no it is not, don't let old algebra notation confuse you d/dx is not a fraction in calculus it is an operator, and the x^n has nothing to do with x raised to the n power how about if I say this let's say y is a function of x